# Discrete Aware Matrix Completion via Convexized $\ell_{0}$-Norm Approximation

### v1

@@ II-A Classical Formulation of the Matrix Completion Problem
s.t. $\quad\text{rank}({\boldsymbol{X}})\leq\mu.$ (6b)
@@ II-B Soft-impute Approach
$\boldsymbol{X}_{t}=\text{SVT}_{\lambda}(\boldsymbol{X}_{t}+P_{\Omega}(\boldsymbol{O}-\boldsymbol{X}_{t})),$ (7)
@@ II-C Discrete-Aware Matrix Completion
with $0\leq p$ and $\text{vec}_{\bar{\Omega}}(\boldsymbol{X})$ represents a vectorization of the matrix $\boldsymbol{X}$, where the entries are chosen corresponding to the given index set $\bar{\Omega}$, being the complementary set to the previously presented set $\Omega$.
@@ III-A Discrete-Aware Regularizer with $\ell_{0}$-Norm Approximation
$r(\boldsymbol{X}|0)$ $=\sum_{k=1}^{|\mathcal{A}|}||\text{vec}_{\bar{\Omega}}(\boldsymbol{X})-a_{k}\boldsymbol{1}||_{0}=\sum_{k=1}^{|\mathcal{A}|}\sum_{j=1}^{|\mathcal{X}|}||\boldsymbol{x}_{j}-a_{k}||_{0}$
$\approx\sum_{k=1}^{|\mathcal{A}|}\bigl{(}\underbrace{|\mathcal{X}|}_{const.}-\sum_{j=1}^{|\mathcal{X}|}\frac{\alpha}{|\boldsymbol{x}_{j}-a_{k}|^{2}+\alpha}\bigr{)}$
$=-\sum_{k=1}^{|\mathcal{A}|}\sum_{j=1}^{|\mathcal{X}|}\frac{\alpha}{|\boldsymbol{x}_{j}-a_{k}|^{2}+\alpha},$ (21)
where $|\mathcal{X}|$ denotes the number of elements in the set obtained after applying $\text{vec}_{\bar{\Omega}}(\cdot)$ to $\boldsymbol{X}$. Additionally, since $|\mathcal{X}|$ is a constant term, it can be neglected in the following due to working on an optimization problem.
@@ III-B Reformulation via Fractional Programming
$-\sum_{k=1}^{|\mathcal{A}|}\sum_{j=1}^{|\mathcal{X}|}\frac{\alpha}{|\boldsymbol{x}_{j}-a_{k}|^{2}+\alpha}=\sum_{k=1}^{|\mathcal{A}|}\sum_{j=1}^{|\mathcal{X}|}\beta_{k,j}^{2}|\boldsymbol{x}_{j}-a_{k}|^{2},$ (23)
@@
$\boldsymbol{b}=\sum_{k=1}^{|\mathcal{A}|}a_{k}[\beta_{k,1}^{2},\beta_{k,2}^{2},\cdots\beta_{k,|\mathcal{X}|}^{2}]^{\intercal},$ (24a)
$\boldsymbol{B}=\sum_{k=1}^{|\mathcal{A}|}\text{diag}(\beta_{k,1}^{2},\beta_{k,2}^{2},\cdots\beta_{k,|\mathcal{X}|}^{2})\succ 0,$ (24b)
@@ IV Numerical Results
$\text{NMSE}\triangleq\frac{||P_{\Omega^{c}}(\boldsymbol{X}-\boldsymbol{O})||^{2}_{F}}{||P_{\Omega^{c}}(\boldsymbol{O})||^{2}_{F}}.$ (32)

### v2

@@ II-A Classical Formulation of the Matrix Completion Problem
s.t. $\quad\text{rank}({\boldsymbol{X}})\leq\nu.$ (6b)
@@ II-B Soft-impute Approach
$\boldsymbol{X}_{t}=\text{SVT}_{\lambda}(\boldsymbol{X}_{t-1}+P_{\Omega}(\boldsymbol{O}-\boldsymbol{X}_{t-1})),$ (7)
@@ II-C Discrete-Aware Matrix Completion
with $0\leq p$, $|\mathcal{A}|$ being the cardinality of the set and $\text{vec}_{\bar{\Omega}}(\boldsymbol{X})$ representing a vectorization of the matrix $\boldsymbol{X}$, where the entries are chosen corresponding to the given index set $\bar{\Omega}$, being the complementary set to the previously presented set $\Omega$.
@@ III-A Discrete-Aware Regularizer with $\ell_{0}$-Norm Approximation
$r(\boldsymbol{X}|0)$ $=\sum_{k=1}^{|\mathcal{A}|}||\text{vec}_{\bar{\Omega}}(\boldsymbol{X})-a_{k}\boldsymbol{1}||_{0}=\sum_{k=1}^{|\mathcal{A}|}\sum_{j=1}^{|\bar{\Omega}|}||\boldsymbol{x}_{j}-a_{k}||_{0}$
$\approx\sum_{k=1}^{|\mathcal{A}|}\bigl{(}\underbrace{|\bar{\Omega}|}_{const.}-\sum_{j=1}^{|\bar{\Omega}|}\frac{\alpha}{|\boldsymbol{x}_{j}-a_{k}|^{2}+\alpha}\bigr{)}$
$=-\sum_{k=1}^{|\mathcal{A}|}\sum_{j=1}^{|\bar{\Omega}|}\frac{\alpha}{|\boldsymbol{x}_{j}-a_{k}|^{2}+\alpha},$ (21)
where $|\bar{\Omega}|$ denotes the number of elements in the set $\bar{\Omega}$. Additionally, since $|\bar{\Omega}|$ is a constant term, it can be neglected in the following due to working on an optimization problem.
@@ III-B Reformulation via Fractional Programming
$-\sum_{k=1}^{|\mathcal{A}|}\sum_{j=1}^{|\bar{\Omega}|}\frac{\alpha}{|\boldsymbol{x}_{j}-a_{k}|^{2}+\alpha}=\sum_{k=1}^{|\mathcal{A}|}\sum_{j=1}^{|\bar{\Omega}|}\beta_{k,j}^{2}|\boldsymbol{x}_{j}-a_{k}|^{2},$ (23)
@@
$\boldsymbol{b}=\sum_{k=1}^{|\mathcal{A}|}a_{k}[\beta_{k,1}^{2},\beta_{k,2}^{2},\cdots\beta_{k,|\bar{\Omega}|}^{2}]^{\intercal},$ (24a)
$\boldsymbol{B}=\sum_{k=1}^{|\mathcal{A}|}\text{diag}(\beta_{k,1}^{2},\beta_{k,2}^{2},\cdots\beta_{k,|\bar{\Omega}|}^{2})\succ 0,$ (24b)
@@ IV Numerical Results
$\text{NMSE}\triangleq\frac{||P_{\bar{\Omega}}(\boldsymbol{X}-\boldsymbol{O})||^{2}_{F}}{||P_{\bar{\Omega}}(\boldsymbol{O})||^{2}_{F}}.$ (32)